OpenStudy (anonymous):

whats the general answer of (x^2)y'+3xy=sinx/x ?

OpenStudy (jamesj):

What's the solution of \[ x^2 y' + 3x y = 0 \]

OpenStudy (jamesj):

I would think about the integrating factor. Then your problem becomes easy.

myininaya (myininaya):

These questions are fun because they linear differential equations :)

OpenStudy (anonymous):

isnt it seperably?

myininaya (myininaya):

The equation James wrote is.

OpenStudy (jamesj):

Integrating factor

myininaya (myininaya):

But the equation you wrote is not.

myininaya (myininaya):

Do you want me to show you the general way to solve these types if it is okay with James?

OpenStudy (anonymous):

yes.

OpenStudy (jamesj):

sure

myininaya (myininaya):

Ok so we have something that can be written in this form: \[y'+p(x)y=q(x)\] Now if we multiply both sides by v>0 such that v'=vp since product rule is vy'+v'y=(vy)' \[vy'+vpy=vq\] So vp=v' so this means we have \[vy'+v'y=vq => (vy)'=vq\] Now we can integrate both sides: \[vy=\int\limits_{}^{}vq dx +C\] And solve for y \[y=\frac{1}{v} \int\limits_{}^{}vq dx +C\] But what is v? Recall v'=vp We can solve this for v by using separation of variables v'=dv/dx by the way and p is a function of x \[\frac{dv}{dx}=vp => \frac{dv}{v}=p dx => \ln(v)=\int\limits_{}^{}p dx+c => v=e^{\int\limits_{}^{} p dx+c} \text{ note take c=0}\] \[ => v=e^{\int\limits_{}^{}p dx} \text{ this is call the integrating factor} \]

myininaya (myininaya):

mistake

myininaya (myininaya):

\[y=\frac{1}{v} \int\limits\limits_{}^{}vq dx +\frac{C}{v}\]

myininaya (myininaya):

i forgot to divide the constant by v above

OpenStudy (anonymous):

for particular solution to this one, you might be to do variation of parameter

myininaya (myininaya):

Do you have any questions put into this form y'+py=q and then find what we called the integrating factor and if you want you can use that formula I have for finding y

OpenStudy (anonymous):

im trying to solve it but im confused

OpenStudy (anonymous):

it takes a little time for me ^.^

myininaya (myininaya):

I understand work takes time

OpenStudy (anonymous):

see im try but it seems im really dumb first i divide both side by x^2 and get y' + 3y/x = sinx/x^3 second i integrating factor is e^(integral of p) (and p is equal to 3/x right?) and i get e^3lnx equation now looks like y'(e^3lnx) + (e^3lnx)*3y/x = (e^3lnx)*sinx/x^3 what should i do with this 1. its bigger than its initial form /cry

myininaya (myininaya):

or x^3 right?

myininaya (myininaya):

e^(3lnx)=e^(lnx^3)=x^3

myininaya (myininaya):

\[x^3y'+x^3 \cdot \frac{3}{x}y=\sin(x)\] \[x^3y'+3x^2y=\sin(x)\] remember product rule \[(x^3y)'=\sin(x)\]

myininaya (myininaya):

can you get it form there just integrate both sides now

myininaya (myininaya):

from*

myininaya (myininaya):

lol sorry

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

so in the end answer is (c-cosx)/x^3 but im having trouble in typing in this form because my mainboard is in rma.(my current system is amd 3500+ /cry). thnx for help.

OpenStudy (anonymous):

8 hours till TEST

myininaya (myininaya):

lets check I'm going to integrate the last equation I wrote. \[x^3y=-\cos(x)+C\] yes thats right! :) gj

myininaya (myininaya):

i don't know what all that other you are talking about is sorry

myininaya (myininaya):

some kind of computer talk i believe i guess